OpenStudy (anonymous):

Use the quadratic formula to solve the equation 3x^2-5x=1

OpenStudy (anonymous):

This is the quadratic equation (the square root sign should stretch over the -4ac sign too): \[x = \frac{-b ± \sqrt b^2 - 4ac}{2a}\] and you can find your a, b, and c from this formula: \[ ax^{2} + bx + c = 0\] then substitute the values in and solve and i'll check your answer.

OpenStudy (anonymous):

A = 3, B = -5, C = -1

OpenStudy (anonymous):

x= -(-5)^2+/-sqr(-5)^2-4.3.(-1)/2(3) =x=5+/-sqr37/6 Solution set is {5+sqr37/6 or 5-sqr37/6} P/s check

OpenStudy (diyadiya):

Yes ,YOu're right :)